NEED EXCEL FORMULAS FOR ANSWER NOT JUST FINAL NUMBER

Question: ABC HAS A 20 YEAR BOND WITH 7% COUPON PAID ANNUALLY AND IS PRICED $944.00.  IF ISSUANCE COSTS ARE 3%, WHAT IS COST OF BOND TO ABC?
Answer #1

This can be solved using financial calcultor or excel
Using financial calculator
N=20
PMT=7%*1000=70
PV=-944*(1-3%)=-915.68
FV=1000
CPT I/Y=7.85%

using excel
=RATE(20,70,-915.68,1000)
=7.85%